In the discrete probability distribution what is the sum of all probabilities?
The sum should equal to 1.
What are the mean and the standard deviation of a proportion?
The mean of a proportion, p, is X/n; where X is the number of instances & n is the sample size; and its standard deviation is sqrt[p(1-p)]
What is an example of a variable that is not normally distributed?
A variable that shows serious departure from the classic bell-shaped, or "Gaussian", curve is described as being not normally distributed. This departure could take the form of skew and/or kurtosis and/or multi modality.
An example might be weekly wages. If you drew a histogram of a population's earnings you would most likely see a distribution skewed significantly toward the right. That is, toward the higher incomes.
Another example is height. If you drew a histogram of a population's height you would see a bimodal distribution. One peak for males and another peak for females. The distribution of height for males and females might be normal when looked at individually, but not normal when you combine them.
The probability is one in fifty-two.
When a single card is drawn from an ordinary 52 card deck find the probability of getting a black 7?
P(7 of Clubs) or P(7 of Spades) = 1/52 + 1/52 = 0.0385
or 1/26 chance
What shape would the probability distribution have for completely certain returns?
a straight, vertical line, i.e., zero variability
A random variable is a function that assigns unique numerical values to all possible outcomes of a random experiment.
A real valued function defined on a sample space of an experiment is also called random variable.
How many possible outcomes of tossing a coin 3 times have exactly 2 heads?
Three outcomes are possible: THH or HTH or HHT, so the probability of getting exactly 2 heads from three tosses of a fair coin is 3x0.53=0.375
In general, the number of different ways of getting k heads from n coin tosses is given by n! / ( k! x (n-k)! ), which is known as the binomial coefficient. The "!" notation is "the factorial"; for some number s, s! = s x (s-1) x (s-2) ... 2 x 1, so that 3! = 3x2x1, 5! = 5x4x3x2x1 and so on.
Thus your question can be answered by replacing n and k in the formula above, such that the number of ways is 3! / ( 2! x 1!) = 3. This is true for all experiments where you perform some experiment with two possible outcomes n times and record how many times one of the events occurs. Another typical example would be the number of times you roll a 4 after 10 rolls of a dice, and there are 10! / 6!x4! = 210 ways of doing this.
The first M&M picked is not restricted to one colour, so the probability of it being an acceptable M&M (one whose colour is not already taken) is 1. For the second M&M, there are 15 left in the bag and 3 of these are not acceptable - because we just chose the other one of that colour. So the probability of the second being acceptable is 12/15 = 0.8. For the third, we are down to only 8 acceptable M&Ms (two colours to choose from, four of each) but have 14 left in the bag (two have been removed and presumably eaten already). This is 8/14 = 4/7 ~ 0.57. For the last pick, there are only four acceptable M&Ms left in the bag of 13; 4/13 is about 0.31 or 31%.
Multiplying these together, we see that the probability of having one of each of four colours is equal to 0.14065934065934065934065934065934 or 0.141 to 3 places. This is the same as 14.1% in total.
Stockbrokers and traders are human beings. They may read an event differently, have different predictions on it's consequences, or react to it differently based on emotions, media attention, or something else entirely.
Off the top of my head, a perfect F-ratio would be 1.00 which is never possible. All F-ratios will be greater than one so the numerator has to be greater than denominator.
This is called a random sample.
Measurements that fall beyond three standard deviations?
It is one of the informal definitions for an outlier.
How many permutations are possible of the word yard?
Imagine you have four empty buckets in which to put any of the four letters Y, A, R, D into. At first you have four letters that can be placed in any of the four empty buckets. Once you've placed a letter in a bucket you only have three letters and three empty buckets to choose from. And so on... So there are 4x3x2x1 = 24 permutations of the word YARD.
What is the measures that fall beyond three standard deviations of the mean called?
You may be referring to the statistical term 'outlier(s)'. Also, there is a rule in statistics called the '68-95-99 Rule'. It states that in a normally distributed dataset approximately 68% of the observations will be within plus/minus one standard deviation of the mean, 95% within plus/minus two standard deviations, and 99% within plus/minus three standard deviations. So if your data follow the classic bell-shaped curve, roughly 1% of the measures should fall beyond three standard deviations of the mean.
How is basketball related to math?
wow great question, well for sure that i know basketball has alot to do with math, well, it has everything to do with math. the angles and the height of basketball player the strength used to throw the ball and pretty much everything else. Also, it has to do a to with angles
1. Compute the square of the difference between each value and the sample mean.
2. Add those values up.
3. Divide the sum by n-1. This is called the variance.
4. Take the square root to obtain the Standard Deviation.
Why divide by n-1 rather than n in the third step above?
In step 1, you compute the difference between each value and the mean of those values. You don't know the true mean of the population; all you know is the mean of your sample. Except for the rare cases where the sample mean happens to equal the population mean, the data will be closer to the sample mean than it will be to the true population mean.
The value you compute in step 2 will probably be a bit smaller (and can't be larger) than what it would be if you used the true population mean in step 1. To make up for this, divide by n-1 rather than n.
But why n-1?
If you knew the sample mean, and all but one of the values, you could calculate what that last value must be. Statisticians say there are n-1 degrees of freedom.
---- == ==
What is the difference between a normal distribution and a standard normal distribution?
when you doesnt have information about the real mean of a population and use the estimation of mean instead of the real mean , usually you use t distribution instead of normal distribution.
* * * * *
Intersting but nothing to do with the question!
If a random variable X is distributed Normally with mean m and standard deviation s, then Z = (X-m)/s has a standard Normal distribution.
Z has mean 0 and standard deviation = 1 (or Variance = sd2 = 1).
What are some causes of skewed data distributions?
The cause of skewed data distributions are extreme values, also know as outliers. For example imagine taking the weights of people you see on the street. If you have 9 cheerleaders' weights and then the weight of a sumo wrestler mixed into the averages this skews the data. This makes the mean much higher because of the one extreme value. Instead of the data being distributed normally, it is distributed with a positive skew. If there is a really small extreme value instead of a really large one, then the data has a negative skew. This could be the heights of people on the street, but one of them would be a midget. The mean is made lower by that one extreme value.
Perhaps, little person is a more politically correct term in our day.
Why should data be normally distributed in order to calculate probabilities using z score tables?
whatis normal distribution
What is the probability of a boys basketball making it into the hoop versus a girls ball?
well, i play basket ball and to tell you the truth, it's 50 50 on that. But that's if the players have had the same amount of experience. Boys and girls are both great at sinking' hoops.
I would expect it to land on a number of 4 or higher 50% of the time or roughly 24 times.